# On Neural Networks as Infinite Tree-Structured Probabilistic Graphical Models

### v3

@@ 6 Limitations and Future Work
Limitations of the present work and directions for future work include establishing formal results about how closely batch- and layer-normalization approximate Markov network normalization when using non-sigmoid activations, establishing theoretical results relating HMC in the neural network to Gibbs sampling in the large treewidth-1 Markov network, and obtaining empirical results for HMC with non-sigmoid activations. Also of great interest is comparing HMC and other PGM algorithms to Shapley values, Integrated Gradients, and other approaches for assessing the relationship of some latent variables to each other or to inputs and/or outputs in a neural network. Finally, the large treewidth-1 PGM is a substantial approximation to the direct PGM of a DNN. After training the DNN and hence the large treewidth-1 model, can we fine-tune with a less-approximate approach, perhaps based on loopy belief propagation or other approximate algorithms often used in PGMs?
@@ Acknowledgements
The authors would like to thank Sayan Mukherjee, Samuel I. Berchuck, Youngsoo Baek, Andrew Allen, William H. Majoros, David Carlson, Juan Restrepo and Houssam Nassif for their helpful discussion about the theoretical work. We are also grateful to Mengyue Han and Jinyi Zhou for their technical support.
@@
Appendix A Proof of Theorem 1: Bayesian Belief Net and Markov Net Equivalence
Theorem.
@@ Appendix A Proof of Theorem 1: Bayesian Belief Net and Markov Net Equivalence
We use $DOM(\phi)$ to designate the variables in a potential $\phi$. Because the nodes and structures of $M$ and $N$ agree, we will refer to the parents, children, ancestors, and descendants of any node in $M$ to designate the corresponding nodes in $N$. Likewise we will refer to the input and output variables of $M$ as those nodes of $N$ that have no parents and no children, respectively. Because $M$ has treewidth 1, each node of $M$ d-separates its set of ancestors from its set of descendants and indeed from all other nodes in $M$. As a result, it is known that the partition function can be computed efficiently in treewidth-1 Markov networks, for example by the following recursive procedure $f$ defined below. Let $V_{0}$ be the empty set of variables, and let $V_{1}$ be the input variables of $M$. Let $Ch(V)$ denote the children of any set $V$ of variables in $M$, and similaly let $Pa(V)$ denote the parents of $V$. For convenience, when $V$ is a singleton we drop the set notation and let $V$ denote the variable itself. For all natural numbers $i\geq 0$:

### v5

@@ 6 Conclusion, Limitations, and Future Work
Limitations of the present work and directions for future work include establishing formal results about how closely batch- and layer-normalization can be modified to approximate Markov network normalization when using non-sigmoid activations, establishing theoretical results relating HMC in the neural network to Gibbs sampling in the large treewidth-1 Markov network, and obtaining empirical results for HMC with non-sigmoid activations. Also of great interest is comparing HMC and other PGM algorithms to Shapley values, Integrated Gradients, and other approaches for assessing the relationship of some latent variables to each other or to inputs and/or outputs in a neural network. We note that the large treewidth-1 PGM is a substantial approximation to the direct PGM of a DNN – in other words, the PGM whose structure exactly matches that of the DNN. In future work, we will explore other DNN fine-tuning methods, perhaps based on loopy belief propagation or other approximate algorithms often used in PGMs, that may allow us to more closely approximate inference in this direct PGM.
Another direction for further work is in the original motivation for this work. Both DNNs and PGMs are often used to model different components of very large systems, such as the entire gene regulatory network in humans. For example, in the National Human Genome Research Institute (NHGRI) program Impact of Genetic Variation on Function (IGVF), different groups are building models of different parts of gene regulation, from genotypic variants or CRISPRi perturbations of the genome, to resulting changes in transcription factor binding or chromatin remodeling, to post-translational modifications, all the way to phenotypes characterized by changes in the expression of genes in other parts of the genome IGVF Consortium (2024). Some of these component models are DNNs and others are PGMs. As a community we know from years of experience with PGMs that passing the outputs of one model to the inputs of another model is typically less effective than concatenating them into a larger model and fine-tuning and using this resulting model. But this concatenation and fine-tuning and usage could not be done with a mixture of PGM and DNN components until now. Having an understanding of DNN components as PGMs enables their combination with PGM components, and then performing fine-tuning and inference in the larger models using algorithms such as the new HMC algorithm theoretically justified, developed, and then evaluated in this paper. Furthermore, the same HMC approach can be employed to reason just as easily from desired gene expression changes at the output nodes back to variants or perturbations at the input nodes that are predictive of the desired changes. Ordinarily, to reason in reverse in this way in a DNN would require special invertible architectures or training of DNNs that operate only in the other direction such as diffusion. Experiments evaluating all these uses of HMC (or other approximate algorithms in PGMs such as loopy belief propagation or other message passing methods) are left for future work.
@@ Acknowledgements
The authors would like to thank Sayan Mukherjee, Samuel I. Berchuck, Youngsoo Baek, David B. Dunson, Andrew S. Allen, William H. Majoros, Jude W. Shavlik, Sriraam Natarajan, David E. Carlson, Kouros Owzar, and Juan Restrepo for their helpful discussion about the theoretical work. We are also grateful to Mengyue Han and Jinyi Zhou for their technical support.
@@
Appendix A Bayesian Belief Net and Markov Net Equivalence
We don’t claim the following theorem is new, but we provide a proof because it captures several components of common knowledge to which we couldn’t find a single reference.
Theorem A.1.
@@ Appendix A Bayesian Belief Net and Markov Net Equivalence
We use $DOM(\phi)$ to designate the variables in a potential $\phi$. Because the nodes and structures of $M$ and $N$ agree, we will refer to the parents, children, ancestors, and descendants of any node in $M$ to designate the corresponding nodes in $N$. Likewise we will refer to the input and output variables of $M$ as those nodes of $N$ that have no parents and no children, respectively. Because $M$ has treewidth 1, each node of $M$ d-separates its set of ancestors from its set of descendants and indeed from all other nodes in $M$. As a result, it is known that the partition function can be computed efficiently in treewidth-1 Markov networks, for example by the following recursive procedure $f$ defined below. Let $V_{0}$ be the empty set of variables, and let $V_{1}$ be the input variables of $M$. Let $Ch(V)$ denote the children of any set $V$ of variables in $M$, and similarly let $Pa(V)$ denote the parents of $V$. For convenience, when $V$ is a singleton we drop the set notation and let $V$ denote the variable itself. For all natural numbers $i\geq 0$: